Assume Company XYZ’s currently has capital Structure C under Method 1 where the WACC is 13.20% based on an effective tax rate of 30% and a cost of debt of 10.43%. There are 200 shares currently outstanding. The company just reported (T=0) FCFF of 300 and FCFE of 250. All cash flows are expected to grow at 20% for each of the next three years, then at 6% beginning in year four (T=4) in perpetuity (or indefinitely). If you were an activist investor trying to gain control of Company XYZ, what is the most you would pay per share for the hostile takeover?

A. $6.97

B. $9.62

C. $11.55

D. $15.49

E. $31.36
Answer #1

Answer: Option D:$15.49

For Capital Structure C Debt: Equity =1:1

So debt weightage =0.5

and Equity weightage r=0.5

WACC=debt weightage*(1-Tax rate)*Cost of debt+Cost of equity*Equity weightage

13.2%=0.5*(1-30%)*10.43%+0.5*r

r=19.0%

given FCFE for T=0 is 250 and it is increasing for 20% for 3 years and 6% for perpetuity

cash flow for year 1 = 250*(1+20%)=300

PV of perpetuity cash flow (T=3) =Cash flow for year 3*(1+6%)/(r-6%)=432*(1+6%)/(19%-6%)=3522.46

PV of FCFE= FCFE/(1+r)^n

n= year in which cash flow occur

for n=1

PV of FCFE= 300/(1+19%)^1= 252.10

Year 0 1 2 3
cash flow 250 300 360 432
PV of Cash flow for Perpetuity 3522.46
FCFE 250 300 360 3954.46
PV of FCFE 250.00 252.10 254.22 2346.64
Total sum of PV of FCFE 3102.96

So share price for hostile take over=3102.96/200=15.49